OpenStudy (anonymous):

Which number cannot be used in place of x in the following rational expression? x+6/11x+1

OpenStudy (watchmath):

The denominatro can't be zero. It is zero when 11x+1=0, i.e., when x=-1/11. So we can't use x=-1/11.

OpenStudy (anonymous):

which number cannot be used in place of x in the following rational expression x+7/x(2)-64 we cannot use x= or x=

OpenStudy (watchmath):

Give it a try!. For which x we have \(x^2-64=0\) ?